OpenStudy (anonymous):

Help me please and thank you! . Choose the symbol that will create a true sentence when x equals 6. 42 + (x – 3)2 __ 28 (Points : 1) = ≥ > <

OpenStudy (anonymous):

its c! :)

OpenStudy (anonymous):

\(42 + (x-3)^2 -- 28 \) Right?

OpenStudy (anonymous):

Put \(x = 6\)

OpenStudy (anonymous):

\(42 + (6-3)^2 -- \ 28\)

OpenStudy (anonymous):

What is \(6-3\) = ??

OpenStudy (anonymous):

its 3

OpenStudy (anonymous):

\(42 + (3)^2 -- \ 28\)

OpenStudy (anonymous):

what is \(3^2\) = ??

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Oops..!!

OpenStudy (anonymous):

\(3 \times 2 = 6\) \(3^2 = 3 \times 3 = ??\)

OpenStudy (anonymous):

Power means to multiply the number that many times like: \(4^3\), it means to multiply \(4\), \(3\) times, ie \(= 4 \times 4\times 4 = 64\)

OpenStudy (anonymous):

oh okay i messed up my bad.

OpenStudy (anonymous):

So, \(3^2 = ??\)

OpenStudy (anonymous):

\(3^2\) means to multiply \(3\), \(2\) times ie \(= 3 \times 3 = ??\)

OpenStudy (anonymous):

9 right?

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

\(42 + 9 - 28\)

OpenStudy (anonymous):

\(42+9 =??\)

OpenStudy (anonymous):

i dont think the question is asking for a power by stating (x-3)2, i think its asking basically (6)-3x2

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

so solve the equation by substituting 6 for x and then see which symbol makes the statement true. if that makes sense :/

OpenStudy (anonymous):

To me, the question is: